# Efficient generative adversarial networks using linear additive-attention Transformers

### v1

@@ Abstract
Although the capacity of deep generative models for image generation, such as Diffusion Models (DMs) and Generative Adversarial Networks (GANs), has dramatically improved in recent years, much of their success can be attributed to computationally expensive architectures. This has limited their adoption and use to research laboratories and companies with large resources, while significantly raising the carbon footprint for training, fine-tuning, and inference. In this work, we present LadaGAN, an efficient generative adversarial network that is built upon a novel Transformer block named Ladaformer. The main component of this block is a linear additive-attention mechanism that computes a single attention vector per head instead of the quadratic dot-product attention. We employ Ladaformer in both the generator and discriminator, which reduces the computational complexity and overcomes the training instabilities often associated with Transformer GANs. LadaGAN consistently outperforms existing convolutional and Transformer GANs on benchmark datasets at different resolutions while being significantly more efficient. Moreover, LadaGAN shows competitive performance compared to state-of-the-art multi-step generative models (e.g. DMs) using orders of magnitude less computational resources.
@@ I Introduction
This paper presents LadaGAN, an efficient GAN architecture for image generation that is based on linear additive-attention Transformers, which we call Ladaformer. This LadaGAN block is inspired by Fastformer [15], an architecture originally proposed for text processing. We employ Ladaformer in both the generator and the discriminator of LadaGAN, allowing efficient processing of long sequences in both networks. In the generator, this block progressively generates a global image structure from the latent space using attention maps. In the discriminator, the Ladaformer generates attention maps to distinguish real and fake images. Notably, the design of LadaGAN reduces the computational complexity and overcomes the training instabilities often associated with Transformer GANs.
@@ III-A Linear additive attention (Lada)
LadaGAN attention mechanism is heavily inspired by Fastformer’s [15] additive attention  [^1] Not to be confused with Bahdanau’s attention [34]. This efficient $O(N)$ Transformer architecture was originally designed for text processing, achieving comparable long-text modeling performance to the original dot-product attention at a fraction of the computational cost. Instead of computing the pairwise interactions among the input sequence vectors, Fastformer’s additive attention creates a global vector summarizing the entire sequence using a single attention vector computed from the queries.
@@ III-B Ladaformer
Note that this is slightly different from ViTGAN’s self-modulated layer normalization, which injects a vector w computed by passing the latent vector ${\mathbf{z}}$ through a projection network; in contrast, LadaGAN injects ${\mathbf{z}}$ directly. In addition, unlike ViT, ViTGAN, and Fastformer, the LadaGAN generator does not have the residual connection from the output of the attention module to the output of the multi-layer perceptron (MLP).
@@
Figure 2: Local Embedding Expansion. The output of the first Transformer consists of 4 feature maps that are expanded using pixel shuffle (a) to generate a single feature map. Then, the output map from pixel shuffle passes through a convolution (b) to expand to 4 channels. These new 4 feature maps are the input to the second Transformer. In this way, even though the Transformers process sequences of different lengths, the dimensions of the embeddings are independent of the Pixel Shuffle operation.
@@ III-C Generator
LadaGAN generator employs the pixel shuffle operation to progressively transform the latent vector into an image. This operation is a common technique to increase spatial resolution, in which the input is reshaped from $(B,C\times r^{2},H,W)$ to $(B,C,H\times r,W\times r)$, where $r$ is a scaling factor, $B$ is the batch size, $C$ is the number of channels of the output and $H$ and $W$ are the height and width of the input. Although this technique was originally proposed as an efficient alternative to standard ConvNet-based upsampling in super-resolution architectures, it has been widely adopted in image generation Transformers, including recent Transformer GANs. Since a pixel shuffle operation reduces the number of channels $C$ in the input (process (a) in Figure  2), we apply a convolutional layer after such operation to increase the number of channels; we denote this operation as Local Expansion of the Embedding
@@
Given the latent vector ${\mathbf{z}}\in{\mathbb{R}}^{D_{z}}$, and $L$ Transformer blocks, LadaGAN generator operates as follows:
@@ III-E Loss function
LadaGAN employs standard non-saturating logistic GAN loss with R1 gradient penalty [39]. The R1 term penalizes the gradient on real data, allowing the model to converge to a good solution. For this reason, it has been widely adopted in state-of-the-art GANs with convolutional discriminators. More specifically, the loss function is defined as follows:
$\begin{split}\mathcal{L}_{D}=&\mathbb{E}_{{\mathbf{z}}\sim p_{z}({\mathbf{z}})}[\log(D(G({\mathbf{z}})))]\\ &-\mathbb{E}_{{\mathbf{x}}\sim p_{\text{r}}({\mathbf{x}})}[\log(1-D({\mathbf{x}}))]\\ &+\gamma\cdot\mathbb{E}_{{\mathbf{x}}\sim p_{r}({\mathbf{x}})}[(\|\nabla_{{\mathbf{x}}}D({\mathbf{x}})\|^{2}_{2})],\end{split}$ (14)
$\mathcal{L}_{G}=-\mathbb{E}_{{\mathbf{z}}\sim p_{z}({\mathbf{z}})}[\log(D(G({\mathbf{z}})))].$ (15)
Note that the Lada stability allows training $O(N)$ Transformer discriminators with gradient penalties. This is particularly important since dot-product Transformer discriminators and gradient penalties have been found to be incompatible [7], typically causing training instabilities.
@@ IV Experiments
To demonstrate LadaGAN stability, efficiency and competitive performance, we conduct an ablation study to assess the impact of the residual connections, convolutions, and modulation using different $O(N)$ Transformers in the generator, as well as the behavior of a Lada discriminator  [^2] The source code is available at https//github.com/milmor/LadaGAN.. We also evaluate the efficiency of LadaGAN in terms of training data requirements. Finally, we compare the performance and computational complexity of LadaGAN with state-of-the-art single-step and multi-step image generation models.
TABLE I: FID and number of FLOPS for Linformer, Swin-Transformer, Fastformer, Ladaformer, and Lada-Swin generators on CIFAR-10 ($32\times 32$).
Attention Convolutions Residual-MLP Modulation FLOPs FID normal-$\downarrow$
@@
6.59
5.91
6.36
4.88
6.31
6.95
@@ IV-A Experiment setup
To assess the performance of image generation models, we adopt the Fréchet Inception Distance (FID) [43]. In this metric, the distance between visual features of the real data distribution $p_{r}(\cdot)$ and the generated data distribution $p_{g}(\cdot)$ is computed, where the visual features are obtained by encoding the images with a pre-trained Inception-v3 network [44]. Since FID is sensitive to the size of both the real data and the generated data, we follow the same evaluation methodology as previous works for the sake of comparison. Specifically, similar to ViTGAN [7], we compute the FID between all the training images and 50k generated images for CIFAR-10 and between the test images and 19,962 generated images for CelebA. Given the large number of training images in LSUN Bedroom, we calculate the FID between 50k randomly sampled training images and 50k generated images, as done by Ho et al. [4]. Finally, like ADM-IP [6], we compute the FID between 50k randomly selected training images and the complete training set for FFHQ. In addition, we evaluate the complexity of the models in terms of FLOPs, parameters, throughput, and images observed during training.
@@ IV-B Implementation details
We train all models with R1 regularization [39] and the Adam optimizer [45] with $\beta_{1}=0.5$ and $\beta_{2}=0.99$. For all resolutions, the generator learning rate is $0.0002$. For the convolutional discriminators, the learning rate is $0.0004$, while for the Transformer discriminators, we set it to $0.0002$. We use convolutional discriminators for all experiments in subsection IV-C. The initial Ladaformer block generates $8\times 8$ maps with dimension $1024$, followed by a $16\times 16$ Ladaformer with dimension $256$. The last Ladaformer generates maps of $32\times 32$ with dimension $64$. For CIFAR-10, we use pixel-level generation. For CelebA we use patch generation. On the other hand, for FFHQ and LSUN bedroom we stack a convolutional decoder with upsampling in the final Ladaformer block instead of performing patch or pixel generation. The number of heads is $4$, and the MLP dimension is $512$ in all Ladaformers. For CIFAR-10 and FFHQ we use Translation, Color, and Cutout data augmentation [46], and a balanced consistency regularization (bCR) [47] with $\lambda_{real}=\lambda_{fake}=1.0$ and $0.1$ respectively. For CelebA and LSUN bedroom we use Translation and Color data augmentation and do not employ bCR since we do not observe a performance gain.
Figure 5: (a) FFHQ generated images ($128\times 128$). Additive attention maps (b) $32\times 32$, (c) $16\times 16$ and (d) $8\times 8$.
@@ IV-C Ablation studies
We evaluate the image generation quality, efficiency and stability of a Ladaformer generator and compare it with Linformer low-rank attention, Swin-Transformer down-sampling attention, and Fastformer original additive attention. In addition, we evaluate Ladaformer with a Swin-style down-sampling technique, which we call Lada-Swin. We carry out an ablation study to analyze the stability and compatibility of generators based on these attention mechanisms with a convolutional layer and the residual connection of the MLP. In addition, we study the impact of the self-modulated layer normalization on the Ladaformer performance. To make the models comparable, we use the same training configuration (see IV-B) and set some of the hyperparameters so that all generators take approximately the same number of FLOPs. In particular, all Fastformer hyperparameters are the same as Ladaformer, whereas for Linformer, we use a $k=64$ and for Swin Transformer and Lada-Swin, a window size of $8\times 8$. For all configurations, we employ a convolutional discriminator identical to FastGAN discriminator but without batch normalization. The main block of this architecture consists of a convolutional residual block with an average pool residual connection (similar to the green blocks in Figure 4).
@@
Dataset Discriminator D learning rate LEE G Embedding size G-FLOPs D-FLOPs FID normal-$\downarrow$
@@
CelebA Convolutional 0.0004 {1024, 256, 64} 0.7B 0.7B 3.36
Convolutional 0.0004 {1024, 256, 128} 0.9B 0.7B 3.28
Lada 0.0002 {1024, 256, 64} 0.7B 0.9B 2.91
Lada 0.0002 {1024, 256, 128} 0.9B 0.9B 3.12
LSUN Bedroom Convolutional 0.0004 {1024, 256, 64} 0.9B 0.9B 9.81
Convolutional 0.0004 {1024, 512, 256} 4.3B 0.9B 6.14
Lada 0.0002 {1024, 256, 64} 0.9B 1.1B 6.59
Lada 0.0002 {1024, 512, 256} 4.3B 1.1B 4.90
@@
Table I shows the FID scores and number of FLOPs for all the evaluated attention mechanisms and configurations. As can be observed, the Ladaformer with a convolutional layer and without the residual connection achieves the lowest FID with a similar number of FLOPs. In general, adding a convolutional layer has a positive effect on Ladaformer, a negative effect for Linformer and practically no effect for Fastformer. This shows not only that the combination of Ladaformer and convolutions does not lead to training instabilities, but also that it can provide noticeable benefits for the quality of the generated images. This is because the locality of the convolutional layer might complement the long-range dependencies of the additive attention map that is propagated by the element-wise operation in equation 3. Note that, as opposed to Lada, Fastformer attention mechanism compresses the representation by computing a second additive attention map for the keys instead of propagating it through an element-wise operation, which seems to prevent the benefits of the convolutional layer.
Interestingly, Fastformer and Ladaformer obtained slightly lower FIDs without the residual connection. However, this only occurs when the Transformer is modulated; without SLN, it has the opposite effect on FID. On the other hand, the performance of Linformer, which is built upon dot-product attention, deteriorates considerably when removing the residual connection. This suggests that Transformers based on linear additive attention mechanisms are less dependent on such shortcuts to propagate the gradients properly and that SLN might be playing a similar role in these kinds of connections. However, this behavior indicates that more investigation is required into the residual connections and modulation of Transformers that do not employ dot-product attention.
@@
The effect of the residual connection on the gradients of Linformer and Fastformer can be seen in Figure  5. Removing the residual connection considerably increases the gradient norms of Linformer in both the generator and discriminator, while the gradients in the generator become widely unstable. In contrast, although removing the residual connection in LadaGAN leads to slightly higher FIDs and larger gradient norms, in general they remain stable. This is in part because of the R1 penalty and its compatibility with LadaGAN, as discussed in subsection  IV-C. Note that these effects are not observed when adding or removing the convolutional layer (see Figure  5).

### v4

@@ Abstract
Although the capacity of deep generative models for image generation, such as Diffusion Models (DMs) and Generative Adversarial Networks (GANs), has dramatically improved in recent years, much of their success can be attributed to computationally expensive architectures. This has limited their adoption and use to research laboratories and companies with large resources, while significantly raising the carbon footprint for training, fine-tuning, and inference. In this work, we present a novel GAN architecture which we call LadaGAN. This architecture is based on a linear attention Transformer block named Ladaformer. The main component of this block is a linear additive-attention mechanism that computes a single attention vector per head instead of the quadratic dot-product attention. We employ Ladaformer in both the generator and discriminator, which reduces the computational complexity and overcomes the training instabilities often associated with Transformer GANs. LadaGAN consistently outperforms existing convolutional and Transformer GANs on benchmark datasets at different resolutions while being significantly more efficient. Moreover, LadaGAN shows competitive performance compared to state-of-the-art multi-step generative models (e.g. DMs) using orders of magnitude less computational resources [^1] The source code is available at https://github.com/milmor/LadaGAN.
@@ I Introduction
This paper presents LadaGAN, a novel efficient GAN architecture for image generation that is based on a linear additive-attention Transformer, which we call Ladaformer. We employ Ladaformer in both the generator and the discriminator of LadaGAN, allowing efficient processing of long sequences in both networks. In the generator, this block progressively generates a global image structure from the latent space using attention maps. In the discriminator, the Ladaformer generates attention maps to distinguish real and fake images. Notably, the design of LadaGAN reduces the computational complexity and overcomes the training instabilities often associated with Transformer GANs.
@@ III Method
In this section, we introduce the proposed LadaGAN architecture. The key component of LadaGAN is a linear additive attention Transformer which is combined with convolutional layers to build the generator and discriminator blocks. To the best of our knowledge, this is the first GAN architecture that uses linear additive attention and convolutional layers in both the generator and the discriminator. Note that the design of a GAN architecture with a Transformer discriminator has proven to be challenging due to the computing cost of the dot-product attention and the training instabilities associated with the gradient penalty [7] commonly used in GANs.
@@ III-A Linear additive attention (Lada)
LadaGAN attention mechanism is inspired by Fastformer’s [15] additive attention  [^2] Not to be confused with Bahdanau’s attention [34]. This efficient $O(N)$ Transformer architecture was originally designed for text processing, achieving comparable long-text modeling performance to the original dot-product attention at a fraction of the computational cost. Instead of computing the pairwise interactions among the input sequence vectors, Fastformer’s additive attention creates a global vector summarizing the entire sequence using a single attention vector computed from the queries.
@@ III-B Ladaformer
Note that this is slightly different from ViTGAN’s self-modulated layer normalization, which injects a vector ${\mathbf{w}}$ computed by passing the latent vector ${\mathbf{z}}$ through a projection network; in contrast, LadaGAN injects ${\mathbf{z}}$ directly. In addition, unlike ViT, ViTGAN, and Fastformer, the LadaGAN generator does not have the residual connection from the output of the attention module to the output of the multi-layer perceptron (MLP).
@@
Figure 3: Local Embedding Expansion. The output of the first Transformer consists of 4 feature maps that are expanded using pixel shuffle (a) to generate a single feature map. Then, the output map from pixel shuffle passes through a convolutional layer (b) to expand to 4 channels. These new 4 feature maps are the input to the second Transformer. In this way, even though the Transformers process sequences of different lengths, the dimensions of the embeddings are independent of the Pixel Shuffle operation.
@@ III-C Generator
LadaGAN generator employs the pixel shuffle operation to progressively transform the latent vector into an image. This operation is a common technique to increase spatial resolution, in which the input is reshaped from $(B,C\times r^{2},H,W)$ to $(B,C,H\times r,W\times r)$, where $r$ is a scaling factor, $B$ is the batch size, $C$ is the number of channels of the output and $H$ and $W$ are the height and width of the input. Although this technique was originally proposed as an efficient alternative to standard ConvNet-based upsampling in super-resolution architectures, it has been widely adopted in image generation Transformers, including recent Transformer GANs. Since a pixel shuffle operation reduces the number of channels $C$ in the input (process (a) in Figure  3), we apply a convolutional layer after such operation to increase the number of channels; we denote this operation as Local Expansion of the Embedding
@@
Given the latent vector ${\mathbf{z}}\in{\mathbb{R}}^{D_{{\mathbf{z}}}}$, and $L$ Transformer blocks, LadaGAN generator operates as follows:
@@ III-E Loss function
LadaGAN employs standard non-saturating logistic GAN loss with $R_{1}$ gradient penalty [39]. The $R_{1}$ term penalizes the gradient on real data, allowing the model to converge to a good solution. For this reason, it has been widely adopted in state-of-the-art GANs with convolutional discriminators. More specifically, the loss function is defined as follows:
$\begin{split}\mathcal{L}_{D}=&-\mathbb{E}_{{\mathbf{x}}\sim P_{{\mathbf{x}}}}[\log(D({\mathbf{x}}))]\\ &-\mathbb{E}_{{\mathbf{z}}\sim P_{{\mathbf{z}}}}[1-\log(D(G({\mathbf{z}})))]\\ &+\gamma\cdot\mathbb{E}_{{\mathbf{x}}\sim P_{{\mathbf{x}}}}[(\|\nabla_{{\mathbf{x}}}D({\mathbf{x}})\|^{2}_{2})],\end{split}$ (14)
$\begin{split}\mathcal{L}_{G}=&-\mathbb{E}_{{\mathbf{z}}\sim P_{{\mathbf{z}}}}[\log(D(G({\mathbf{z}})))].\end{split}$ (15)
@@ IV Experiments
TABLE I: FID and number of FLOPs for the LadaGAN generator using Linformer, Swin-Transformer, Fastformer, Ladaformer, and Lada-Swin attention mechanisms on CIFAR-10 ($32\times 32$).
Attention Convolutions Residual-MLP Modulation FLOPs FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ Prec $\uparrow$ Rec $\uparrow$
@@
6.47
5.88
6.29
4.82
6.19
5.87
5.76
4.98
5.43
4.57
6.28
4.73
9.24
9.18
9.26
9.69
9.35
9.33
0.56
0.58
0.57
0.58
0.57
0.57
0.58
0.56
0.57
0.59
0.57
0.53
To demonstrate LadaGAN stability, efficiency and competitive performance, we conduct an ablation study to assess the impact of the residual connections, convolutions, and modulation using different $O(N)$ Transformers in the generator, as well as the behavior of a Lada discriminator. We also evaluate the efficiency of LadaGAN in terms of training data requirements. Finally, we compare the performance and computational complexity of LadaGAN with state-of-the-art single-step and multi-step image generation models.
@@ IV-A Experiment setup
Figure 5: Gradient magnitudes over all parameters of the LadaGAN generator and discriminator, and FID evaluation of the attention mechanisms (a), convolutional layer (b), and MLP residual connection (c).
To assess the performance of image generation models, we adopt the Fréchet Inception Distance (FID) [43]. In this metric, the distance between visual features of the real data distribution and the generated data distribution is computed, where the visual features are obtained by encoding the images with a pre-trained Inception-v3 network [44]. Unlike FID, the spatial FID (sFID) captures spatial relationships by employing spatial features rather than standard pooled features. In addition to FID and sFID, we also report Precision and Inception Score (IS) to measure the fidelity of the generated samples and Recall to measure their diversity. Since FID is sensitive to the size of both the real data and the generated data, we follow the same evaluation methodology as previous works for the sake of comparison. Specifically, similar to ViTGAN [7], we compute the FID between all the training images and 50k generated images for CIFAR-10 and between the test images and 19,962 generated images for CelebA. Given the large number of training images in LSUN Bedroom, we calculate the FID between 50k randomly sampled training images and 50k generated images, as done by Ho et al. [4]. Finally, like ADM-IP [6], we compute the FID between 50k randomly selected training images and the complete training set for FFHQ. Additionally, we evaluate the complexity of the models in terms of FLOPs, parameters, throughput, and images observed during training.
@@ IV-B Implementation details
We train all models with $R_{1}$ regularization [39] and the Adam optimizer [45] with $\beta_{1}=0.5$ and $\beta_{2}=0.99$. For all resolutions, the generator learning rate is $0.0002$. For the convolutional discriminators, the learning rate is $0.0004$, while for the Transformer discriminators, we set it to $0.0002$. We use convolutional discriminators for all experiments in subsection IV-C. The initial Ladaformer block generates $8\times 8$ maps with dimension $1024$, followed by a $16\times 16$ Ladaformer with dimension $256$. The last Ladaformer generates maps of $32\times 32$ with dimension $64$. For CIFAR-10, we use pixel-level generation. For CelebA we use patch generation. On the other hand, for FFHQ and LSUN bedroom we stack a convolutional decoder with upsampling in the final Ladaformer block instead of performing patch or pixel generation. The number of heads is $4$, and the MLP dimension is $512$ in all Ladaformers. For CIFAR-10 and FFHQ we use Translation, Color, and Cutout data augmentation [46], and a balanced consistency regularization (bCR) [47] with $\lambda_{real}=\lambda_{fake}=1.0$ and $0.1$ respectively. For CelebA and LSUN bedroom we use Translation and Color data augmentation and do not employ bCR since we do not observe a performance gain.
@@ IV-C Ablation studies
We evaluate the image generation quality, efficiency and stability of a Ladaformer generator and compare it with Linformer low-rank attention, Swin-Transformer down-sampling attention, and Fastformer original additive attention. In addition, we evaluate Ladaformer with a Swin-style down-sampling technique, which we call Lada-Swin. We carry out an ablation study to analyze the stability and compatibility of generators based on these attention mechanisms with a convolutional layer and the residual connection of the MLP. We also examine the impact of the self-modulated layer normalization on the Ladaformer performance. To make the models comparable, we use the same training configuration (see IV-B) and set some of the hyperparameters so that all generators take approximately the same number of FLOPs. In particular, all Fastformer hyperparameters are the same as Ladaformer, whereas for Linformer, we use a $k=64$ and for Swin Transformer and Lada-Swin, a window size of $8\times 8$. For all configurations, we employ a convolutional discriminator identical to FastGAN discriminator but without batch normalization. The main block of this architecture consists of a convolutional residual block with an average pool residual connection (similar to the green blocks in Figure 4).
@@
Dataset D-type D-lr LEE G-emb sizes G-FLOPs D-FLOPs FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ Prec $\uparrow$ Rec $\uparrow$
@@
CelebA Conv 0.0004 {1024, 256, 64} 0.7B 0.7B 3.43 11.98 3.31 0.67 0.55
Conv 0.0004 {1024, 256, 128} 0.9B 0.7B 3.36 11.79 3.36 0.66 0.58
Lada 0.0002 {1024, 256, 64} 0.7B 0.9B 2.89 11.24 3.26 0.67 0.58
Lada 0.0002 {1024, 256, 128} 0.9B 0.9B 3.04 11.16 3.26 0.67 0.59
[…]
Conv 0.0004 {1024, 512, 256} 4.3B 0.9B 5.82 15.17 2.25 0.49 0.36
Lada 0.0002 {1024, 256, 64} 0.9B 1.1B 6.46 20.35 2.14 0.45 0.38
Lada 0.0002 {1024, 512, 256} 4.3B 1.1B 4.60 14.81 2.22 0.52 0.41
  • $\ast$
    Convolutional decoder with nearest neighbor upsampling instead of patch generation.
@@
Table I shows the FID scores, IS Precision, Recall, and number of FLOPs for all the evaluated attention mechanisms and configurations. As can be observed, the Ladaformer with a convolutional layer and without the residual connection achieves the best evaluation with a similar number of FLOPs. In general, adding a convolutional layer has a positive effect on Ladaformer, a negative effect for Linformer and practically no effect for Fastformer. This shows not only that the combination of Ladaformer and convolutions does not lead to training instabilities, but also that it can provide noticeable benefits for the quality of the generated images. This is because the locality of the convolutional layer might complement the long-range dependencies of the additive attention map that is propagated by the element-wise operation in equation 3. Note that, as opposed to Lada, Fastformer attention mechanism compresses the representation by computing a second additive attention map for the keys instead of propagating it through an element-wise operation, which seems to prevent the benefits of the convolutional layer.
Interestingly, Fastformer and Ladaformer obtained slightly lower FIDs without the residual connection. However, when Ladaformer employs SLN, the improvement in the FID when removing this connection is stronger. On the other hand, the performance of Linformer, which is built upon dot-product attention, deteriorates considerably when removing the residual connection. This suggests that Transformers based on linear additive attention mechanisms are less dependent on such shortcuts to propagate the gradients properly and that SLN might be playing a similar role in these kinds of connections. However, this behavior indicates that more investigation is required into the residual connections and modulation of Transformers that do not employ dot-product attention.
@@
The effect of the residual connection on the gradients of Linformer and Fastformer can be seen in Figure  5. Removing the residual connection considerably increases the gradient norms of Linformer in both the generator and discriminator, while the gradients in the generator become widely unstable. In contrast, although removing the residual connection in LadaGAN leads to slightly higher FIDs and larger gradient norms, in general they remain stable. This is in part because of the $R_{1}$ penalty and its compatibility with LadaGAN, as discussed in subsection  IV-C. Note that these effects are not observed when adding or removing the convolutional layer (see Figure  5).